# How Linguistics Learned to Stop Worrying and Love the Language Models

### v1

@@ Abstract
Language models can produce fluent, grammatical text. Nonetheless, some maintain that language models don’t really learn language and also that, even if they did, that would not be informative for the study of human learning and processing. On the other side, there have been claims that the success of LMs obviates the need for studying linguistic theory and structure. We argue that both extremes are wrong. LMs can contribute to fundamental questions about linguistic structure, language processing, and learning. They force us to rethink arguments about learning and are informative for major questions in linguistic theory. But they do not replace linguistic structure and theory. We offer an optimistic take on the relationship between language models and linguistics.
@@ 1 Introduction
Fast forward to 2025. Norm has worked for decades on a variety of diverse languages, developing explanatory theories of linguistic structure. Claudette got more and more interested in engineering, amassing huge amounts of data, and training statistical models. Norm and Claudette slowly fell out of touch over the years and haven’t seen each other for a while. They are planning another lunch: what should they talk about?
Just like Claudette always dreamed of, she now has machines that appear to be able to talk with her in English, producing grammatical and seemingly sensible utterances. How relevant is it that the architecture of Claudette’s machines seems to have nothing to do with the structure of language as identified by Norm and other linguists and cognitive scientists? Claudette might not care: she just wanted a system that worked and this stuff works. On the other hand, what if Norm was right about the nature of language—does that mean the machines aren’t actually as impressive as Claudette thinks, because they are relying on shallow pattern matching? Or are Claudette’s machines evidence that Norm’s theories were wrong? More generally, what insights do these machines have for the scientific study of language?
@@
This negative view takes several forms. One argument is that Language Models (yes, Claudette’s machines are Language Models; henceforth LMs) are so fundamentally different in their architecture from people, and have access to so much more data, that whatever they are doing is so different as to be totally irrelevant for humans (Bolhuis et al.,, 2024; Chomsky et al.,, 2023). Another argument is that neural sequence architectures could learn anything, so the fact that they learn language (to the extent that they do) is uninformative (Rawski and Baumont,, 2023; Moro et al.,, 2023; Chomsky,, 2023; Chomsky et al.,, 2023; Bolhuis et al.,, 2024). Under this view, neural networks are like mathematical clay that can be formed into any shape (Robinson,, 2023): the fact that they can be formed into something like the shape of language doesn’t tell us anything about language, any more than the fact that we can form clay into the shape of a human body tells us anything about anatomy. A distinct but related negative approach has been to deny that language models have learned, or ever could learn, the putatively key properties of human language, and thus are either irrelevant to language science or do not challenge any existing theories (Lan et al.,, 2024; Fox and Katzir,, 2024). We believe these negative views are short-sighted: language models do learn nontrivial aspects of linguistic structure, and they do give important insights that change how we should think about language. As language scientists, we ignore them at our peril.
@@
An opposite approach is to dismiss traditional linguistic ideas of structure, on the grounds that they turned out to be either useless or of negative value in engineered language technologies (Piantadosi,, 2023). Under this view, decades of work on the structure of language turned out to be barking up the wrong tree, and the theory of language should be rebuilt on top of the only computational foundation that has produced a machine that can actually use language: large neural networks. We believe this view is also short-sighted. It throws out hard-won analytical discoveries about the structure of language and leaves us adrift in terms of a scientific theory of language, without a clear way to approach the question of why human language is the way it is, or even a sense of what the interesting questions are. Moreover, language models are currently most successful in languages with internet-scale data and are biased towards standardized prestige dialects, neglecting under-resourced languages and dialects (Blasi et al.,, 2022). Work using language models will benefit from linguistic work focused on communities outside this current narrow band—which requires the expertise of documentary linguists, sociolinguists, anthropologists, and community stakeholders.
@@
Thus, we advocate a third view (joining others in cognitive science and philosophy who have advocated for this middle ground, for example, Smolensky,, 1988; Millière,, 2024; Portelance and Jasbi,, 2023; Pater,, 2019; Potts,, 2025): large language models are not a complete theory of language—in fact, no one has such a theory—but they are hugely informative about language and its structure, learning, processing, and relationship with the larger structure of the mind. Language models have set off an intellectual explosion in cognitive science, machine learning, philosophy of mind, and other fields, in which longstanding ideas have been overturned, truly novel ideas are emerging, and disciplinary boundaries are dissolving. Linguistics has a chance to stand at the center of this huge intellectual foment, and would be remiss to isolate itself intellectually on the basis that language models don’t look like any particular existing theory. We believe that language science can contribute (and in fact already has contributed) to the development of language models, and language models can contribute (and in fact already have contributed) insights about language. Norm and Claudette have a lot to talk about.
@@ 2.1 A brief history of statistical language learning
The generative school of modern linguistics arose from an argument against the idea that the structure of language could be learned from the statistics of language data. Chomsky, perhaps most famously illustrated this by contrasting “(1) Colorless green ideas sleep furiously” with “(2) Furiously sleep ideas green colorless.” Even though neither sentence makes semantic sense, (1) is clearly grammatical in a way that (2) isn’t. As Chomsky, (1957) put it: “The notion ‘grammatical in English’ cannot be identified in any way with the notion ‘high order of statistical approximation to English.’ It is fair to assume that neither sentence (1) nor (2) (nor indeed any part of these sentences) has ever occurred in an English discourse. Hence, in any statistical model for grammaticalness, these sentences will be ruled out on identical grounds as equally ‘remote’ from English.”
The conclusion from these arguments was that linguistic structure could only be characterized in terms of formal systems, based on rules or constraints and operating over structured arrays of symbols (Chomsky,, 1965). Early on, the expectation was that such systems would form the basis for language technologies such as machine translation (Hays,, 1960; Hutchins,, 1979). It was believed that these formal systems should be constructed by linguists since the task of learning them from data was hard or impossible. Yet despite myriad efforts to build machine translation systems, grammatical parsers, and other tools, linguistic competence remained elusive for machines. Symbolic approaches that sought to elucidate rules and structures often proved unable to capture all the exceptions and complexity that characterized natural language as it is used.
By the late 1980s and 1990s, statistical learning had a major renaissance in natural language processing (Brown et al.,, 1990; Manning and Schütze,, 1999; Pereira,, 2000), and started to reappear in the human language learning literature as well (Saffran et al.,, 1996). Yet despite success at various NLP tasks, these techniques seemed frustratingly unable to get past approaches that simply counted up words and phrases (Wang and Manning,, 2012; Arora et al.,, 2017). The connectionist movement in the 1980s and 90s seemed promising (Rumelhart and McClelland, 1986b, ; Smolensky,, 1988; Elman, 1990b, ) (and turned out to be the most important forerunner of modern language models), but there were fundamental concerns about the ability of these approaches to scale up and to represent the rules and structures that seem to characterize language (Pinker and Prince,, 1988), despite accounts of how connectionist models could in principle implement rule-like symbolic behavior (Smolensky,, 1990).
@@
Against this backdrop, the fortunes of connectionist models started changing in the 2010s—as a result of new techniques and, most importantly, due to increased computational power that enabled training neural models much more efficiently (Hinton et al.,, 2006). By 2016, neural models showed rudiments of structure-dependent generalizations like subject–verb agreement (Linzen et al.,, 2016). Over the coming years, the success of models at acquiring structure-dependent linguistic abilities continued to grow (Futrell et al., 2019a, ; Wilcox et al.,, 2018; Manning et al.,, 2020; Hu et al.,, 2020; Warstadt and Bowman,, 2022; Mahowald et al.,, 2024).
The growth from early neural models in, e.g., 2011 to now is, from a historical perspective, remarkable and surprising to virtually everyone in the field. Sutskever et al., (2011) introduced an at-the-time state-of-the-art recurrent neural network that produced output like “In the show’s agreement unanimously resurfaced. The wild pasteured with consistent street forests were incorporated by the 15th century BE.” and praised it as a “surprisingly good language model,” noting the “richness of their vocabularies,” that the “text is mostly grammatical,” and “parentheses are usually balanced over many characters.” Compared to today’s language models, the consistent street forests seem a long way away: 2011 might as well have been the 15th century BE. Models like ChatGPT or Claude, based on large neural networks, produce seemingly endless streams of highly fluent and entirely grammatical (if not always perfectly humanlike) text, and form the basis of the most successful NLP approaches.
@@ 2.2 Neural LMs learn nontrivial linguistic structure
The rapid development of these capacities has been treated with healthy skepticism. The field of linguistics has provided a deep body of work exploring the subtleties and complexities of even seemingly simple linguistic utterances, and LMs have access to vast amounts of data. There is good reason to be skeptical and wonder whether the models are simply memorizing, or giving the appearance of linguistic competence without really having it.
@@
The most naïve way that one might think to test for grammatical competence in language models would be to see whether they assign higher probability to grammatical strings than ungrammatical ones. However, this would not work, even for an ideal statistical model, for reasons related to the point that Chomsky, (1957) was making. For example, the ungrammatical string ‘Snails died the old’ has a probability of $\sim 2^{-49}$ under the GPT-2 language model, whereas the grammatical string ‘The ancient crustaceans expired’ has a lower probability of $\sim 2^{-55}$, simply because it uses lower-frequency words (Wilcox et al., 2023a, ).
@@
But the failure of this naïve approach does not mean that langauge models lack grammatical knowledge. In general—for language models and in real language use—there are many factors that influence the probability of a string, of which the grammar of the language (as represented in the language model or in the mind of a human speaker) is only one: word frequency, utterance length, online processing constraints such as memory limitations, and plausibility given world knowledge all feed into the probability of an utterance. Indeed, not only utterance probabilities, but also comprehension accuracy, reaction time, and indeed any psychometric dependent variable are affected by all of these factors jointly—including the subjective grammaticality judgments that form the basis of much of the formal syntax literature (Kluender and Kutas,, 1993; Hofmeister et al.,, 2013; Mahowald et al.,, 2016; Lau et al.,, 2017). Therefore, if we want to look for evidence that language models represent linguistic structure, we need to search more carefully. [^2] Nor is it sensible to evaluate models based on simply prompting them and asking whether sentences are grammatical or not (Hu and Levy,, 2023) (as done by Dentella et al.,, 2023). To see this, consider that many linguists have held that grammatical knowledge can be represented by a formal system like a context-free grammar; even if a context-free grammar is a good model of linguistic knowledge, it could not be prompted in natural language to respond about whether a sentence is grammatical or not. The ability to do explicit grammaticality judgments when prompted in natural language is not a necessary consequence of grammatical knowledge.
@@
The key is to isolate linguistic structure from these other factors through controlled experimental studies and through probing and manipulating LMs’ internals. Experimentally, from sufficient performance data, one may infer an underlying formal cognitive structure, no matter whether the implementation substrate is a brain or a neural network (Piantadosi and Gallistel,, 2024). This is the standard procedure in linguistics, where data consisting primarily of acceptability judgments is used to postulate underlying linguistic competence. This approach can be applied just as well to LMs.
@@
One form of such a study is to conduct behavioral comparisons of minimally different sentences. For example, LMs may be fed sentences such as The keys to the cabinet are on the table and the ungrammatical The keys to the cabinet is on the table, and the conditional probabilities assigned to the verb form are versus is are compared: if the model understands how agreement works in English, then the probability for are should be higher than is. This comparison controls for plausibility, since the context is the same for the two alternatives. The task could not be solved by any $n$-gram model with $n<5$ (and higher-order $n$-gram models can be ruled out by adding more words between keys and is/are in the stimuli). The lexical frequency of is versus are can be controlled through a more elaborate experimental design, with four conditions in a $2\times 2$ design crossing the grammatical number of the subject with the grammatical form of the verb (as done by Marvin and Linzen,, 2018). These are usual procedures in psycholinguistics, where experimenters develop controlled sets of sentences and then measure dependent variables like reading times. The same methodology can be applied to language models with probability as the dependent variable, aiming to ascertain whether models are sensitive to linguistic structure (Linzen et al.,, 2016; Futrell et al., 2019b, ).
@@
Such studies have revealed behavioral patterns consistent with neural networks representing formal linguistic structures, in cases such as subject–verb agreement (Linzen et al.,, 2016; Bernardy and Lappin,, 2017; Gulordava et al.,, 2018), filler–gap dependencies (Wilcox et al.,, 2018; Wilcox et al., 2023a, ; Kobzeva et al.,, 2023; Suijkerbuijk et al.,, 2023), and recursive embedding of clauses (Futrell et al., 2019b, ; Wilcox et al., 2019a, ; Hu et al.,, 2020), all of which involve highly nontrivial formal structures which statistical models failed to capture in previous work. Example results for subject–verb agreement from GPT-2 are shown in Figure 1: we see that grammatical verb forms are relatively more probable than matched ungrammatical verb forms in all but a few cases (human accuracy in producing the right verb forms in such sentences is 85%; Marvin and Linzen,, 2018). [^3] It has been objected that, while humans make errors in verb agreement and other syntactic constructions, they have an underlying error-free competence which may be distinguished from their errorful performance (Chomsky,, 1965, Ch. 1), whereas the neural networks only have errorful performance with no underlying competence (Fox and Katzir,, 2024). We believe this argument is mistaken for two reasons. First, it is indeed possible to define (multiple) competence–performance distinction(s) in neural networks, for example by separating the knowledge contained in embeddings versus the performance of a linear decoder in extracting that knowledge (Pimentel et al.,, 2020; White et al.,, 2021; Csordás et al.,, 2024). Second, even for humans, we only ever have access to performance data, and we infer mental constructs such as competence from it in order to explain that data (Piantadosi and Gallistel,, 2024). We may make the same inferences based on neural network performance data. Competence is not a datum to be explained, but rather a theoretical construct that is useful for explaining performance data—it is a real pattern (Dennett,, 1989; Nefdt,, 2023) in the sense we will discuss in Section 4.1. The results indicate that the model can represent the non-local structural dependency between the subject of the sentence and the matrix verb.

### v2

@@ Abstract
Language models can produce fluent, grammatical text. Nonetheless, some maintain that language models don’t really learn language and also that, even if they did, that would not be informative for the study of human learning and processing. On the other side, there have been claims that the success of LMs obviates the need for studying linguistic theory and structure. We argue that both extremes are wrong. LMs can contribute to fundamental questions about linguistic structure, language processing, and learning. They force us to rethink arguments and ways of thinking that have been foundational in linguistics. While they do not replace linguistic structure and theory, they serve as model systems and working proofs of concept for gradient, usage-based approaches to language. We offer an optimistic take on the relationship between language models and linguistics.
@@ 1 Introduction
Fast forward to 2025. Norm has worked for decades on a variety of diverse languages, developing sophisticated theories of linguistic structure. Claudette got more and more interested in engineering, amassing huge amounts of data, and training statistical models. Norm and Claudette slowly fell out of touch over the years. They are planning a reunion lunch: what should they talk about?
Just like Claudette always dreamed of, she now has machines that can talk with her in English, producing grammatical and seemingly sensible utterances. How relevant is it that the architecture of Claudette’s machines seems to have nothing to do with the structure of language as identified by Norm and other linguists and cognitive scientists? Claudette might not care: she just wanted a system that worked, and this stuff works. On the other hand, what if Norm was right about the nature of language—does that mean the machines aren’t actually as impressive as Claudette thinks, because they are relying on shallow pattern matching? Or are Claudette’s machines evidence that Norm’s theories were wrong? More generally, what insights do these machines have for the scientific study of language?
@@
This negative view takes several forms. One argument is that Language Models (yes, Claudette’s machines are Language Models; henceforth LMs) are so fundamentally different in their architecture from people, and have access to so much more data, that whatever they are doing is so different as to be totally irrelevant for humans (Chomsky et al.,, 2023; Fox and Katzir,, 2024; Bolhuis et al.,, 2024). Some have denied that language models have actually learned, or ever could learn, the putatively key properties of human language, and thus are either irrelevant to language science or do not challenge any existing ways of thinking (Lan et al.,, 2024; Fox and Katzir,, 2024). A different and more subtle argument is that neural network sequence models could learn to approximate anything, so the fact that they seem to learn language (to the extent that they do) is uninformative (Rawski and Baumont,, 2023; Moro et al.,, 2023; Chomsky,, 2023; Chomsky et al.,, 2023; Collins,, 2024; Bolhuis et al.,, 2024). Under this view, LM architectures are like epicycles, the computational technique used by Ptolemy to predict the motions of the planets and the sun in a model that placed the Earth at the center of the universe (de Santillana,, 1955; Flynn,, 2013). The problem with epicycles is that they can approximate any trajectory arbitrarily well (at the cost of great complexity), so the mere fact that they could be used to capture the motions of the planets with high accuracy tells us nothing about the real nature of that motion. Similarly, if neural networks are just general pattern approximators, then the fact that they can approximate language doesn’t tell us anything about language.
We believe these negative views are short-sighted: language models do learn nontrivial aspects of linguistic structure, and they do give important insights that change how we should think about language. As language scientists, we ignore them at our peril.
An opposite approach is to dismiss traditional theories of linguistic structure, on the grounds that they turned out to be either useless or of negative value in developing the only known systems that can actually use language as humans do (Jelinek,, 2004; Piantadosi,, 2023). Under this view, decades of work on the structure of language turned out to be barking up the wrong tree, and the theory of language should be rebuilt on top of the only computational foundation that has been demonstrated to be able to produce and comprehend language to a human level: large neural networks. In our experience, this view is widespread in some engineering and application-focused communities.
@@
This view is also short-sighted. It throws out hard-won analytical discoveries about the structure of language and leaves us adrift in terms of a scientific theory of language, without a clear way to approach the question of why human language is the way it is, or even a sense of what the interesting questions are. Moreover, language models are currently most successful in English and other languages with internet-scale data (Blasi et al.,, 2022). A more complete approach to the science of language will draw on the expertise of documentary linguists, sociolinguists, anthropologists, and community stakeholders, and it will integrate the insights from decades of linguistic inquiry.
@@
Thus, we advocate a third view (joining fellow travelers in linguistics, cognitive science, and philosophy who have advocated for some form of this middle ground, e.g., Smolensky,, 1988; Pater,, 2019; Portelance and Jasbi,, 2023; McGrath et al.,, 2024; Millière,, 2024; Potts,, 2025; Chesi,, 2025): language models are not a complete theory of language—in fact, no one has such a theory—but they are hugely informative about language and its structure, learning, processing, and relationship with the larger structure of the mind. Language models have set off an intellectual explosion in cognitive science, machine learning, philosophy of mind, and other fields, in which longstanding ideas have been overturned; novel ideas are emerging; and disciplinary boundaries are dissolving. Linguistics has a chance to stand at the center of this huge intellectual ferment, and would be remiss to isolate itself intellectually on the basis that language models don’t look like existing theory. Language science can contribute (and in fact already has contributed) to the development of language models, and language models can contribute (and in fact already have contributed) insights about language. Norm and Claudette have a lot to talk about.
@@ 2.1 A brief history of statistical language learning
The generative school of modern linguistics arose from an argument against the idea that the structure of language could be learned from the statistics of language data. Chomsky, perhaps most famously illustrated this by contrasting (1) “Colorless green ideas sleep furiously” with “(2) Furiously sleep ideas green colorless.” Even though neither sentence makes semantic sense, (1) is clearly grammatical in a way that (2) isn’t. As Chomsky, (1957) put it: “The notion ‘grammatical in English’ cannot be identified in any way with the notion ‘high order of statistical approximation to English.’ It is fair to assume that neither sentence (1) nor (2) (nor indeed any part of these sentences) has ever occurred in an English discourse. Hence, in any statistical model for grammaticalness, these sentences will be ruled out on identical grounds as equally ‘remote’ from English.”
The effective conclusion from these arguments was that linguistic structure could only be characterized in terms of formal systems, based on rules or constraints and operating over structured arrays of symbols (Chomsky,, 1965). Early on, the expectation was that such systems would form the basis for language technologies such as machine translation and question answering systems (Hays,, 1960; Winograd,, 1972; Hutchins,, 1981). It was believed that these formal systems should be constructed by linguists since the task of learning them from data was hard or impossible. Yet despite myriad efforts to build machine translation systems, grammatical parsers, and other tools, linguistic competence remained elusive for machines. Symbolic approaches that sought to elucidate rules and structures often proved unable to capture all the exceptions and complexity that characterized natural language as it is used.
By the late 1980s and 1990s, statistical learning had a major renaissance in natural language processing (Brown et al.,, 1990; Manning and Schütze,, 1999; Pereira,, 2000), and started to reappear in the human language learning literature as well (Saffran et al.,, 1996). Yet despite success at various natural language processing tasks, these techniques seemed frustratingly unable to get past approaches that simply counted up words and phrases (Wang and Manning,, 2012; Arora et al.,, 2017). The connectionist movement in the 1980s and 90s seemed promising (Rumelhart and McClelland,, 1986; Smolensky,, 1988; Elman, 1990a, ) (and turned out to be the most important forerunner of modern language models), but there were well-justified concerns about the ability of these approaches to scale up and to represent the rules and structures that seem to characterize language (Pinker and Prince,, 1988), despite accounts of how connectionist models could in principle implement rule-like symbolic behavior (Smolensky,, 1990).
@@
Against this backdrop, the fortunes of connectionist models started changing in the 2010s—as a result of new techniques and, most importantly, due to increased computational power that made training neural models much more efficient (Hinton et al.,, 2006). By 2016, neural models showed rudiments of grammatical generalizations like subject–verb agreement (Linzen et al.,, 2016). Over the coming years, the success of models at acquiring linguistic abilities continued to grow (Futrell et al., 2019a, ; Wilcox et al.,, 2018; Manning et al.,, 2020; Hu et al.,, 2020; Warstadt and Bowman,, 2022; Mahowald et al.,, 2024).
The growth from early neural models in, e.g., 2011 to now is remarkable from a historical perspective, and was surprising to virtually everyone in the field at the time. Sutskever et al., (2011) introduced an at-the-time state-of-the-art recurrent neural network that produced output like “In the show’s agreement unanimously resurfaced. The wild pasteured with consistent street forests were incorporated by the 15th century BE.” and praised it as a “surprisingly good language model,” noting the “richness of their vocabularies,” that the “text is mostly grammatical,” and “parentheses are usually balanced over many characters.” Compared to today’s language models, the consistent street forests seem a long way away: 2011 might as well have been the 15th century BE. Today’s language models (e.g., ChatGPT or Claude) produce seemingly endless streams of highly fluent and grammatical (if not always perfectly humanlike) text and form the basis of the most successful natural language processing approaches.
@@ 2.2 Neural LMs learn nontrivial linguistic structure
The rapid development of these capacities has been treated with healthy skepticism. LMs have access to vast amounts of data and could be simply memorizing or relying on cheap heuristics — giving the appearance of linguistic competence without really having it.
@@
The most naïve way that one might think to test for grammatical competence in language models would be to see whether they assign higher probability to grammatical strings than ungrammatical ones. However, this would not work, even for an ideal statistical model, for reasons related to the point that Chomsky, (1957) was making. For example, the ungrammatical string “Snails died the old” has a probability of $\sim 2^{-49}$ under the GPT-2 language model, whereas the grammatical string “The ancient crustaceans expired” has a lower probability of $\sim 2^{-55}$, simply because it uses lower-frequency words (Wilcox et al., 2023a, ).
@@
But the failure of this naïve comparison does not mean that language models lack grammatical knowledge. In general—for language models and in real language use— many factors influence the probability of a string, of which the grammar of the language (as represented in the language model or in the mind of a human speaker) is only one: word frequency, utterance length, online processing constraints such as memory limitations, and plausibility given world knowledge all feed into the probability of an utterance. Indeed, not only utterance probabilities, but also comprehension accuracy, reaction time, and indeed any psychometric dependent variable are affected by all of these factors jointly—including the subjective grammaticality judgments that form the basis of much of the formal syntax literature (Kluender and Kutas,, 1993; Hofmeister et al.,, 2013; Mahowald et al.,, 2016; Lau et al.,, 2017). Therefore, if we want to look for evidence that language models represent linguistic structure, we need to search more carefully. [^2] Nor is it necessarily sensible to evaluate models based on simply prompting them and asking whether sentences are grammatical or not (Hu and Levy,, 2023) (as done by Dentella et al.,, 2023). To see this, consider that many linguists have held that grammatical knowledge can be represented by a formal system like a context-free grammar; even if a context-free grammar is a good model of linguistic knowledge, it could not be prompted in natural language to respond about whether a sentence is grammatical or not. Grammatical knowledge does not imply that a system can make explicit grammaticality judgments when prompted in natural language.
@@
The key is to isolate linguistic structure from these other factors through controlled experimental studies and through probing LMs’ internal states. Experimentally, from sufficient performance data, one may infer an underlying formal cognitive structure, no matter whether the implementation substrate is a brain or a neural network (Piantadosi and Gallistel,, 2024). This is the standard procedure in linguistics, where data consisting primarily of acceptability judgments is used to postulate underlying linguistic competence. This approach can be applied just as well to LMs.
@@
One form of such a study is to conduct behavioral comparisons of minimally different sentences. For example, LMs may be fed sentences such as “The keys to the cabinet are on the table” and the ungrammatical “The keys to the cabinet is on the table”, and the conditional probabilities assigned to the verb form “are” versus “is” are compared. If the model understands how agreement works in English, then the probability for “are” should be higher than “is”. This comparison controls for plausibility, since the context is the same for the two alternatives. The task could not be solved by any $n$-gram model with $n<5$ (and higher-order $n$-gram models can be ruled out by adding more words between “keys” and “is/are” in the stimuli). The lexical frequency of “is” versus “are” can be controlled through a more elaborate experimental design, with four conditions in a $2\times 2$ design crossing the grammatical number of the subject with the grammatical form of the verb (as done by Marvin and Linzen,, 2018). These are usual procedures in psycholinguistics, where experimenters develop controlled sets of sentences and then measure dependent variables like reading times. The same methodology can be applied to language models with probability as the dependent variable, aiming to ascertain whether models are sensitive to linguistic structure (Linzen et al.,, 2016; Futrell et al., 2019b, ).
@@
Such studies have revealed behavioral patterns consistent with neural networks representing formal linguistic structures, in cases such as subject–verb agreement (Linzen et al.,, 2016; Bernardy and Lappin,, 2017; Gulordava et al.,, 2018), filler–gap dependencies (Wilcox et al.,, 2018; Wilcox et al., 2023a, ; Kobzeva et al.,, 2023; Suijkerbuijk et al.,, 2023), and recursive embedding of clauses (Futrell et al., 2019b, ; Wilcox et al., 2019a, ; Hu et al.,, 2020), all of which involve highly nontrivial formal structures which statistical models failed to capture in previous work. Example results for subject–verb agreement from GPT-2 are shown in Figure 1: we see that grammatical verb forms are relatively more probable than matched ungrammatical verb forms in all but a few cases (human accuracy in producing the right verb forms in such sentences is 85%; Marvin and Linzen,, 2018). [^3] It has been objected that, while humans make errors in verb agreement and other syntactic constructions, they have an underlying error-free competence which may be distinguished from their errorful performance (Chomsky,, 1965, Ch. 1), whereas the neural networks only have errorful performance with no underlying competence (Fox and Katzir,, 2024). We believe this argument is mistaken for two reasons. First, it is indeed possible to define (multiple) competence–performance distinction(s) in neural networks, for example, by separating the knowledge contained in embeddings versus the performance of a linear decoder in extracting that knowledge (Pimentel et al.,, 2020; White et al.,, 2021; Csordás et al.,, 2024). Second, even for humans, we only ever have access to performance data (including subjective grammaticality judgments), and we infer mental constructs such as competence, I-language, etc., on the basis of that data (Piantadosi and Gallistel,, 2024). We may make the same inferences based on neural network performance data. Competence is not a datum to be explained, but rather a theoretical construct that is useful for explaining performance data—it is a real pattern (Dennett,, 1989; Nefdt,, 2023) in the sense we will discuss in Section 4.1. The results indicate that the model can represent the non-local structural dependency between the subject of the sentence and the matrix verb.